OpenStudy (anonymous):

I have a one-half gallon container of lemonade and want to divide it into one cup servings. A cup is one sixteenth of a gallon. How many cups would I need? A. 6 B. 7 C. 8 D. 10

OpenStudy (anonymous):

I think its 10

OpenStudy (sleepyjess):

If a cup is 1/16 of a gallon, how many cups are in 1 gallon?

OpenStudy (sleepyjess):

Welcome to OpenStudy :)

OpenStudy (anonymous):

thank you & 6?

OpenStudy (sleepyjess):

1 cup = 1/16 gallon 2 cups = 2/16 gallon 3 cups = 3/16 gallon etc.

OpenStudy (sleepyjess):

How much of a gallon would 6 cups be?

OpenStudy (anonymous):

6/16

OpenStudy (sleepyjess):

Yes!

OpenStudy (sleepyjess):

So how many cups would a full gallon be (16/16)?

OpenStudy (anonymous):

6

OpenStudy (sleepyjess):

6 would be 6/16, we need 16/16 :)

OpenStudy (anonymous):

im confused

OpenStudy (sleepyjess):

To make a full gallon, we would need 16 cups of lemonade

OpenStudy (sleepyjess):

But we have half a gallon of lemonade, so how many cups would be in HALF a gallon

OpenStudy (anonymous):

10?

OpenStudy (sleepyjess):

What is half of 16?

OpenStudy (anonymous):

8

OpenStudy (sleepyjess):

Yes!

OpenStudy (sleepyjess):

So one-half of a gallon is 8 cups

OpenStudy (anonymous):

thank you

OpenStudy (sleepyjess):

No problem! If you need any help getting around the site, just let me know :)